a playground, which measures 60m by 40m, is to be doubled in area by extending each side an equal amount. By how much should each side be extended?
Answer by mananth(16949)

Area of present ground = 60*40 = 2400 sq.meters
when doubled it becomes 4800 square meters
Let the increase in length and width be x meters
(60+x)(40+x)= 4800
2400 +60x+40x +x^2= 4800
x^2+100x-2400=0
x^2+120x-20x-2400=0
x(x+120)-20(x+120) =0
(x-20)(x+120)=0
x= 20
The length and width must be increased by 20 meters